# Causal Predictive Optimization and Generation for Business AI

### v1

@@ Abstract.
The sales process involves sales functions converting leads or opportunities to customers and selling more products to existing customers. The optimization of the sales process thus is key to success of any B2B business. In this work, we introduce a principled approach to sales optimization and business AI, namely the Causal Predictive Optimization and Generation, which includes three layers: 1) prediction layer with causal ML 2) optimization layer with constraint optimization and contextual bandit 3) serving layer with Generative AI and feedback-loop for system enhancement. We detail the implementation and deployment of the system in LinkedIn, showcasing significant wins over legacy systems and sharing learning and insight broadly applicable to this field.
@@ 1. Introduction
Sales functions play a critical role in success and revenue growth of any Business-to-Business (B2B) and Software-as-a-Service (SaaS) business. Each member of the sales fleet manages anywhere from a few to hundreds of accounts, and their main objective is to convert leads and upsell to existing customers. Sales Reps spend majority of their time figuring out the right account or person to talk to and what to talk about. Determining the optimality of each component with these accounts on hand is a highly time-consuming task. An efficient business AI engine will greatly improve productivity of sales fleet and is key to the optimization of the sale process and success of the business.
More specifically, addressing the following challenges that exist in the sales ecosystem is critical.
  • Inefficient allocation of sales resources - Absence of causal modeling and insights lead to wasted efforts on low impact accounts while missing high-value opportunities.
  • Complex trade-offs in decision-making – Balancing multiple objectives and constraints in sales (e.g., revenue growth, engagement, and resource limits) is difficult without a structured optimization approach.
  • Lack of trust and adaptability in recommendations – Sales teams struggle with black-box like model outputs, and static systems fail to incorporate real-time feedback, reducing long-term effectiveness.
To address these challenges, we propose Casual Predictive Optimization and Generation (CPOG) as an end-to-end framework solution.
@@ 1.1. Related Work
The application of Machine Learning (ML) in sales has advanced considerably, as companies increasingly harness data to optimize decision-making, enhance customer engagement, and streamline operations.
ML models play a crucial role in account prioritization by guiding sales teams toward high-potential prospects based on behavioral patterns and engagement signals. These models also improve sales forecasting by leveraging both structured data  (Rohaan et al., 2022) and textual data  (Elalem et al., 2023), and uplift modeling  (Gutierrez and Gérardy, 2017) enables more accurate predictions. Additionally, customer segmentation through clustering techniques facilitates better sales planning by identifying key audience groups for targeted strategies  (Gautam and Kumar, 2022). B2B customer churn prediction (De Caigny et al., 2021), customer conversion in e-commerce (Gubela et al., 2019).
In recent years, causal machine learning, which combines the traditional causal inference and modern machine learning, enables predictions of individual heterogeneous effect of future potential action (Rubin, 1978)(Kennedy, 2023)(Van der Laan, 2006)(Gutierrez and Gérardy, 2017). Multiple researches shows the great potential of causal ML application to the industry(Shi et al., 2024; Tang et al., 2023). Optimization frameworks transform predictive scores into actionable recommendations and some recent examples in the literature include neural optimization with adaptive heuristics (NOAH) for an intelligent marketing system (Wei et al., 2024) and data-driven sales optimization with regression and chaotic pattern search (Gaddam et al., 2024). In addition, contextual bandit systems have been proven to provide personalized recommendations (Ban et al., 2024).
The rise of explainable AI (XAI) and generative AI (GAI) is transforming how businesses adopt ML models by making complex sales predictions more interpretable  (Bohanec et al., 2017) and interaction of sales engine more seamless. Model-agnostic, rule-based explanations help interpret the results  (Rajapaksha and Bergmeir, 2022), model explanation methods have been shown to provide narratives and explanations for consumers of model outputs.
There is very limited work yet on end-to-end solutions that enable sales functions to harness the power of ML, and provide combination of Prediction, Optimization and Explanation at large scale. Previously mentioned optimization frameworks developed by (Wei et al., 2024) and (Gaddam et al., 2024) cannot be applied to sales account prioritization optimization because they differ in domain ((Wei et al., 2024) was designed for intelligent marketing) and objective ((Gaddam et al., 2024) focused on lead generation). As far as we know, CPOG is the first sales optimization framework focused on account prioritization. The contributions the CPOG framework are discussed in the next section.
@@ 1.2. Our Contribution
To our knowledge, CPOG is the first end-to-end framework solution for sales optimization in business AI. In addition, CPOG is the first deployed sales optimization engine in industry setting that
  • utilize causal machine learning to measure the incremental impact of sales actions in the prediction layer, which avoid the inefficient use of sales resources;
  • has dedicated optimization layer to make trade-off of multiple different objective and constraints in sales process, and recommend sales action that balances exploration and exploitation;
@@ 2. System Overview
CPOG addresses the challenges of the sales ecosystem through a structured AI-driven framework that optimizes sales decisions. The architecture comprises multiple layers: a Data Foundation layer that utilizes engagement, spending, and rep-interaction data; a Prediction Layer that applies causal machine learning models to estimate long- and short-term impact given sales rep intervention; an Optimization Layer that incorporates business constraints and prioritization logic using mixed-integer programming and bandit algorithms; and a Serving Layer that delivers recommendations with human-interpretable explanations and feedback loops.
In addition, the serving layer integrates Generative AI (GAI) to improve user experience and explainability, providing detailed intuitive insights that increase trust in AI-driven recommendations. By incorporating a feedback loop for continuous refinement, the CPOG framework performs as a cohesive human-in-the-loop system, seamlessly aligning with sales teams to optimize decision-making and maximize revenue impact. Overall, this AI-driven system streamlines account prioritization, personalizes recommendations, and improves sales productivity, ultimately driving business growth. In the subsequent sections, we provide a detailed discussion of each layer and its components.
@@ 3. Prediction Layer
The prediction layer predicts a group of performance metrics which sales functions can influence through sales actions. Denoting the feature matrix for customers as $x$, feature matrix of possible sales actions as $x^{a}$ and the performance metrics as $y$, the prediction layer builds a functional mapping:
$y=f(x,x^{a})$
We consider two types of metrics: (1) monetization metrics with longer feedback periods like revenue, customer loyalty and customer LTV, (2) customer engagement metrics like product utilization and product adoption with shorter feedback periods. The goal of the prediction layer is estimate the impact of sales actions on monetization metrics and predict near term customer engagement metrics. The functional mapping $f(.)$ can be learned from historical data on customer outcomes.
@@ 3.1. Monetization Uplift Models
We follow the potential outcome framework (Rubin, 1974) in causal inference where $Y$ denotes the desired outcome of certain monetization metric, $a$ denotes the treatment s.t. $a\in\{1,0\}$ and $x^{U}$ are the pre-treatment variables e.g product utilization, revenue before the treatment etc.
(1) $Y_{a}=f(x^{U}|a)$
We predict $Y_{a}$ using ML models trained on different subsets of the data i.e. $f(.)$ is trained on data from customers in the control or treatment group. The nature of the ML models used depends on the quantity of data - a simple regression model will suffice for smaller datasets, tree-based models for medium sized data and neural networks for large and complex datasets. The uplift from sales outreach is estimated as follows:
@@
Meta-learners (Athey and Imbens, 2016)(Künzel et al., 2019) are discrete treatment CATE estimators that model typically on two or more surfaces $Y_{t}$. Meta-learners provide the flexibility of selecting both the estimation model used at each stage along with cross-validation for more data-adaptive estimations at each stage. (Künzel et al., 2019) coined the terms ”S-learner”, ”T-learner” and ”X-learner” which are some common meta-algorithms for estimating heterogeneous treatment effect.
@@ 4. Optimization Layer
The optimization layer within the system adeptly transforms predictive scores into actionable recommendations for sales representatives, employing a dual-strategy approach that synergizes both constrained (convex) optimization and contextual bandit methodologies. The constrained optimization component is designed to harmonize key business objectives, including maximizing monetization uplift and enhancing engagement metrics, all while adhering to predefined operational constraints. This component methodically determines the optimal pairings of accounts and sales representatives, ensuring alignment with strategic business goals.
In parallel, the contextual bandit component augments this framework by dynamically adjusting the daily actions assigned to each sales representative. This adjustment mechanism is acutely focused on immediate, actionable items tailored to specific business needs, such as boosting engagement, preventing customer churn, and promoting upsells. Together, these components work in concert to refine the system’s response to evolving market conditions and representative interactions.
@@ 4.1.1. Formulating the Constrained Optimization
Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ represent the set of customer accounts and =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{M}\}$ the set of sales representatives. Define the binary decision variable $a_{i,j}\in\{0,1\}$ to indicate whether account $c_{i}$ is assigned to representative $r_{j}$. The objective is to maximize a weighted combination of monetization uplift ($U_{i}$) and engagement improvement ($E_{i}$), where the weights dynamically adjust.
@@ 4.1.3. Recommendation Rules
This problem is formulated as a Mixed Integer Programming (MIP) problem, which is computationally expensive. To improve tractability:
@@
    Relaxation: The binary variable $a_{i,j}$ is relaxed to continuous values in $[0,1]$, transforming the problem into a Linear Programming (LP) problem.
@@
    Solver: The relaxed LP is solved using SCIP (Perron and Furnon, 2024; Bolusani et al., 2024).
  3. (3)
    Post-Processing: Continuous solutions are rounded to binary values, ensuring capacity constraints are respected.
@@ 4.1.4. Implementation Notes
  • Account Matching: $a_{i,j}$ is the decision variable indicating the assignment of accounts $c_{i}$ to sales representatives $r_{j}$.
  • Action Recommendation: Heuristic algorithms are utilized to decide initial action recommendation on the matched account, based on values of $y^{U}_{i}$ and $y^{E}_{i}$. (see Algorithm  1) This serves as a cold start for contextual bandit ranking in section  4.2.
  • Solver Strategy: We relax the MIP and solve the LP via SCIP (Perron and Furnon, 2024; Bolusani et al., 2024) for efficient offline inference. Though MIP is supported, relaxation ensures speed and stability. We plan to explore ML-augmented MIP solvers (recent work by Bertsimas et al. (Bertsimas et al., 2022)) for future low-latency use cases.
@@
4.1.5. Output
The optimization layer outputs the following for each account (Table 1 & 2):
@@ 4.1.5. Output
Account ID Rep ID gRank111gRank :account ranking across all reps based on objective values rRank222rRank :account ranking within the assigned rep based on objective values
@@ 4.2. Contextual Bandit
After constrained optimization, the recommendations are processed by a contextual bandit layer, which directly interacts with the final serving layer. This layer is integral to the system, leveraging recommendations from upstream modules, such as Monetization Upsell, Monetization Churn, and Low Engagement mapped at the sales representative level, to optimize these recommendations further using user actions and finally sending them to the serving layer. By combining contextual features of recommendations with user actions/feedback, the contextual bandit iteratively refines its policy, dynamically balancing exploration (to improve understanding of less tried recommendations) and exploitation (to serve the most effective recommendations).
Each sales representative is assigned a set of accounts determined by the constraint optimization step. The contextual bandit selects an action from the following predefined set:
  • $a^{B}$ (Boost Engagement)
  • $a^{C}$ (Prevent Churn)
  • $a^{U}$ (Promote Upsell)
These actions are encoded as contextual features within the model to refine decision-making. Over time, the bandit learns which action maximizes cumulative engagement and business value.
4.2.1. Reward Mapping and Objective Function
The contextual bandit framework leverages user feedback to iteratively refine action selection. The system assigns a reward to each action based on observed user responses:
  • $\mathcal{A}=\{a^{B},a^{C},a^{U}\}$ is the set of actions.
  • Reward from user feedback is defined as,
@@ 4.2.1. Reward Mapping and Objective Function
The overarching goal of the contextual bandit layer is to identify the optimal policy $\pi^{*}$ that maximizes cumulative rewards over time. Formally, this objective can be expressed as:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=1}^{T}y^{\pi}_{t}(a_{t},x_{t})\right],$
where:
  • $a_{t}\in\mathcal{A}$ is the action taken at time $t$.
  • $x_{t}$ represents contextual features at time $t$.
  • $y^{\pi}_{t}(a_{t},x_{t})$ denotes the reward from user feedback by following an action policy $\pi$.
This formulation ensures that the system progressively converges to a policy that aligns recommendations with user preferences and optimizes engagement over time.
4.2.2. Contextual Feature Representation
$x_{t}=[x^{A}_{t},x^{S}_{t},x^{R}_{t}],$
where:
  • $x^{R}_{t}$ (Recommendation Features): Captures details like type of recommendation and predicted impact.
These features are extracted using:
with one-hot encoded categorical representations:
[…]
The resulting vector $x_{t}$ ensures that the model evaluates each action within a context-sensitive framework.
@@ 4.2.3. Policy Optimization
Policy optimization is achieved by employing Neural Bandit (Dai et al., 2022; Zhou et al., 2020). Each action’s reward distribution is modeled as:
$y_{a}=h(x_{t},a_{t})+\epsilon$
where:
  • $h(x_{t},a_{t})$ is a unkown functional mapping from feature to the expected reward.
  • $\epsilon$ is a sub-Gaussian noise.
We approximate the unknown function $h(\cdot)$ with a neural network $f_{\theta}(\cdot)$. The algorithm iterate over these steps:
@@
where:
  • $\nabla f(\cdot)$ is the gradient of the neural network output with respect to input features.
  • $H^{-1}$ is the inverse Hessian matrix used to capture model uncertainty.
@@ 5.1.1. Formulating the Explanation Layer
5.1.2. Feature Mapping
A manual mapping translates feature names into comprehensible language. (Example: Table 7 in the Appendix):
5.1.3. Threshold Definitions
Thresholds are defined for each feature and model combination to determine when a specific recommendation is triggered. (Example: Table 8 in the Appendix)

### v2

@@ Abstract.
The sales process involves converting leads or opportunities into customers and selling additional products to existing clients. Optimizing this process is therefore key to the success of any B2B business. In this work, we introduce a principled approach to sales optimization and business AI, Causal Predictive Optimization and Generation, which comprises three layers: (1) a prediction layer using causal ML; (2) an optimization layer with constraint optimization and contextual bandits; and (3) a serving layer featuring Generative AI and feedback loop. We detail the implementation and deployment of this system at LinkedIn, and share learnings and insights broadly applicable to the field.
@@ 1. Introduction
Sales functions play a critical role in the success and revenue growth of any Business-to-Business (B2B) or Software-as-a-Service (SaaS) company. Each member of the sales team manages anywhere from a handful to hundreds of accounts, and their primary objective is to convert leads and upsell existing customers. Sales representatives spend the majority of their time identifying the right account or contact to engage, and deciding what to discuss. Determining the optimal approach for each account is highly time-consuming. Across the B2B and SaaS sectors, sales teams commonly encounter these challenges.
  • Inefficient allocation of resources: Sales reps often spend excessive time on low-impact accounts while overlooking high-value opportunities;
  • Complex decision trade-offs: Balancing targets for revenue growth, customer engagement, and capacity constraints without clear guidance slows down decision-making;
  • Low trust and adaptability in AI-driven recommendations: Opaque, rigid systems impede adoption, and static models rarely incorporate real-time feedback, undermining long-term effectiveness.
An efficient business AI engine can greatly improve sales-team productivity and is key to optimizing the sales process and driving business success. In this paper, we propose Casual Predictive Optimization and Generation (CPOG) as an end-to-end framework solution.
@@ 1.1. Related Work
Machine learning (ML) models play a crucial role in account prioritization by guiding sales teams toward high-potential prospects based on behavioral patterns and engagement signals. These models also improve sales forecasting by leveraging both structured data (Rohaan et al., 2022) and textual data (Elalem et al., 2023), and customer segmentation through clustering techniques facilitates better sales planning by identifying key audience groups (Gautam and Kumar, 2022). B2B customer churn prediction (De Caigny et al., 2021) and e-commerce customer conversion models (Gubela et al., 2019) further demonstrate the value of predictive analytics in sales. However, most existing prediction models do not exploit causal ML methods (Rubin, 1978; Kennedy, 2023; Van der Laan, 2006; Gutierrez and Gérardy, 2017) for estimating the individual heterogeneous effects of potential sales actions, despite promising industry applications in other areas (Shi et al., 2024; Tang et al., 2023).
Separately, optimization frameworks have been developed to turn predictive scores into actionable recommendations—examples include neural optimization with adaptive heuristics (NOAH) for intelligent marketing systems (Wei et al., 2024) and regression and chaotic pattern search (RCPS) for lead generation (Gaddam et al., 2024). Contextual bandit approaches have also shown value for personalized recommendations (Ban et al., 2024). Yet these methods were not designed for end-to-end sales optimization and no existing optimization framework seamlessly integrates into the sales process.
Meanwhile, advances in explainable AI (XAI) and generative AI (GAI) are beginning to transform how businesses consume ML outputs. Model-agnostic, rule-based explanations improve interpretability (Bohanec et al., 2017; Rajapaksha and Bergmeir, 2022), and GAI techniques promise more natural interactions with sales engines.
Despite these advances, there is currently no comprehensive solution that unifies causal prediction, constrained optimization, explainability, and generative serving at enterprise scale for B2B and SaaS sales. The only two frameworks that (remotely) resemble our approach are NOAH (Wei et al., 2024) and RCPS (Gaddam et al., 2024). NOAH targets intelligent marketing rather than sales optimization and does not include explainable AI or generative serving. RCPS addresses only lead generation as a sub-process of sales optimization, and does not incorporate causal ML, multi-objective optimization, XAI, or GAI, nor has it been demonstrated at industry scale.
@@ 1.2. Our Contribution
To our knowledge, CPOG is the first end-to-end framework for sales optimization in business AI. Moreover, it is the first sales optimization engine deployed in an industry setting that:
  • Utilizes causal machine learning to measure the incremental impact of sales actions in the prediction layer, avoiding inefficient use of sales resources;
  • Features a dedicated optimization layer to trade off multiple objectives and constraints in the sales process, and recommend actions that balance exploration and exploitation;
@@ 2. System Overview
CPOG tackles sales-ecosystem challenges with a structured, AI-driven architecture comprising three complementary layers: the Prediction Layer, which applies causal machine-learning models to estimate both short- and long-term impacts of potential sales-rep interventions, enabling more efficient resource allocation and clearer ROI insights; the Optimization Layer, which integrates business constraints and prioritization logic via mixed-integer programming and contextual bandit algorithms to recommend actions that balance multiple objectives (e.g., revenue growth, engagement) while managing exploration–exploitation trade-offs; and the Serving Layer, which delivers these recommendations with human-interpretable explanations, embedded Generative AI components for natural interactions, and a continuous feedback loop that refines the system over time.
Together, these layers form a cohesive human-in-the-loop system that streamlines account prioritization, personalizes outreach, and maximizes revenue impact. In the following sections, we provide a detailed discussion of each layer and its components.
@@ 3. Prediction Layer
The prediction layer predicts a group of performance metrics which sales functions can influence through sales actions. Denoting the feature for customers as $x$, feature of possible sales actions as $a$ and the performance metrics as $y$, the prediction layer builds a functional mapping:
$y=f(x,a)$
We consider two types of metrics: (i) monetization metrics with longer feedback periods like revenue, customer loyalty and customer LTV, (ii) customer engagement metrics like product utilization and product adoption with shorter feedback periods. The goal of the prediction layer is to estimate the impact of sales actions on monetization metrics and predict near term customer engagement metrics. The functional mapping $f(.)$ can be learned from historical data on customer outcomes.
@@ 3.1. Monetization Uplift Models
Let $y$ denote the desired outcome of certain monetization metric, $a$ the sales outreach action, and $x^{U}$ the pre-treatment variables e.g product utilization, revenue before the treatment etc. We predict $y$ using ML models, i.e. $y_{a}=f(x,a)$. The nature of the ML models used depends on the quantity of data: a simple regression model will suffice for smaller datasets, tree-based models for medium sized data and neural networks for large and complex datasets. The uplift from sales outreach when $a\in\{0,1\}$ can estimated as follows:
@@
Besides above ”two” learner (T-learner) formulation, we can also adopt other Meta-learners(Athey and Imbens, 2016)(Künzel et al., 2019) methods such as S-learner, X-learner or DR-learner. Further information are in Appendix A.1.
@@ 4. Optimization Layer
The optimization layer transforms predictive scores into actionable sales recommendations using a dual strategy that combines constrained optimization with contextual bandit methods. The constrained optimization component balances key business objectives, such as maximizing revenue uplift and engagement, against predefined operational constraints, systematically determining the optimal account–representative assignments to align with strategic goals. Simultaneously, the contextual bandit component adapts daily action recommendations in real time, focusing on immediate priorities like boosting engagement, preventing churn, and driving upsells. Together, these techniques enable the system to respond dynamically to evolving market conditions and sales-rep interactions.
@@ 4.1.1. Formulating the Constrained Optimization
Let $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$ represent the set of customer accounts and =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{M}\}$ the set of sales representatives. Define the binary decision variable $a_{i,j}\in\{0,1\}$ to indicate whether account $c_{i}$ is assigned to representative $r_{j}$. The objective is to maximize a weighted combination of monetization uplift ($U_{i}$) and engagement improvement ($E_{i}$).
@@ 4.1.3. Recommendation Rules
The original problem is formulated as a Mixed Integer Programming (MIP) problem. We ”relax” the problem to Linear Programming by changing binary deterministic decision variable $a_{ij}\in\{0,1\}$ to a probability of taking action, i.e., $0\leq a_{ij}\leq 1$, for two reason:
@@
    Ranking: Besides matching of accounts $c$ to reps $r$, we also have needs to form a ranking of account to inform reps the priority on the set of matched account.
@@
    Computation: Directly solving MIP can be computational expensive and are more likely to have convergence problem. Relaxed LP has advantage on computation efficiency and pipeline stability.
We recommend the accounts to reps with following 2 steps:
  • Account Matching and Ranking: Matching is done by rounding $a_{ij}$, i.e., accounts $c_{i}$ is matched to sales reps $r_{j}$ if $\lfloor a_{ij}+0.5\rfloor=1$. Among the set of matched accounts, we rank them by $a_{ij}$. See example results in Table 1.
  • Action Recommendation: Heuristic algorithms are utilized to decide initial action recommendation on the matched account, based on values of $y^{U}_{i}$ and $y^{E}_{i}$. (see Algorithm  1 and example result in Table 2). This serves as a cold start for contextual bandit ranking in section  4.2.
@@ 4.1.4. Implementation Notes
  • Solver Strategy: LP can be solved via SCIP (Perron and Furnon, 2024; Bolusani et al., 2024) in offline manner for typical sale optimization scale. Dualip (Basu et al., 2020) or ML-augmented solvers (Bertsimas et al., 2022) can be used for large scale or low-latency use cases.
@@
Account ID Rep ID gRank111gRank :account ranking across all reps based on $a_{ij}$ rRank222rRank :account ranking within the assigned rep based on $a_{ij}$
@@ 4.2. Contextual Bandit
After constrained optimization, the recommendations are processed by a contextual bandit layer, which decide final action recommendation among Boost_Engagement, Prevent_Churn and Promote_Upsell. The contextual bandit iteratively refines its policy, dynamically balancing exploration (to improve understanding of less tried recommendations) and exploitation (to serve the most effective recommendations).
4.2.1. Model Set-up
Each sales representative is assigned a set of accounts determined by the constraint optimization step. The contextual bandit selects an action $a_{t}\in\mathcal{A}$ to maximize cumulative reward over time:
$\max_{a_{t}\in\mathcal{A}_{t}}\mathbb{E}\left[\sum_{t=1}^{T}y(a_{t},x_{t})\right],$
where, $\mathcal{A}_{t}=\{a^{B},a^{C},a^{U}\}$, $a^{B}$ index Boost_Engagement, $a^{C}$ Prevent_Churn and $a^{U}$ Promote_Upsell; $x_{t}$ represents contextual features at time $t$, and $y(a_{t},x_{t})$ denotes the reward from user feedback, with reward defined as
@@ 4.2.2. Policy Optimization
Policy optimization is achieved by employing Neural Bandit (Dai et al., 2022; Zhou et al., 2020). Each action’s reward distribution is modeled as $y_{a}=h(x_{t},a_{t})+\epsilon$, where $h(x_{t},a_{t})$ is a unkown functional mapping from feature to the expected reward, $\epsilon$ is a sub-Gaussian noise.
In Neural Bandit, we approximate the unknown function $h(\cdot)$ with a neural network $f_{\theta}(\cdot)$. The algorithm iterate over these steps:
@@
where $\nabla f(\cdot)$ is the gradient of the neural network output with respect to input features, $H^{-1}$ is the inverse Hessian matrix used to capture model uncertainty.
@@
4.2.3. Contextual Feature Representation
$x_{t}=[x^{A}_{t},x^{S}_{t},x^{R}_{t}],$
where:
These features are extracted using:
with one-hot encoded categorical representations:
[…]
@@ 5.1.1. Formulating the Explanation Layer
Examples for feature mapping and threshold definition can be found in Appendix Table 5 and Table 6.